### v1

# Pooling Liquidity Pools in AMMs

Abstract
Market fragmentation across multiple Automated Market Makers (AMMs) creates inefficiencies such as costly arbitrage, unnecessarily high slippage and delayed incorporation of new information into prices. These inefficiencies raise trading costs, reduce liquidity provider profits, and degrade overall market efficiency. To address these issues, we propose a modification of the Constant Product Market Maker (CPMM) pricing mechanism, called the Global Market Maker (GMM), which aggregates liquidity information from all AMMs to mitigate these inefficiencies. Through theoretical and numerical analyses, we demonstrate that the GMM enhances profits for both AMMs and traders by eliminating arbitrage opportunities. Additionally, it reduces the profitability of sandwich attacks and minimizes impermanent losses
Keywords: Decentralized Exchanges (DEXs) Automated Market Makers (AMMs) Constant Product Market Maker (CPMM) Arbitrage Slippage Sandwich Attacks Impermanent Loss.
1 Introduction
As cryptocurrentcies’ popularity grows, so it does the need to exchange them for other cryptocurrencies and for fiat currencies. Crypto exchanges have emerged to satisfy these needs. A crypto exchange is a platform that allows users to buy, sell, and trade cryptocurrencies. Crypto exchanges can be centralized or decentralized. Centralized EXchanges (CEXs) are owned and operated by a single entity, which holds users’ funds and executes trades. Decentralized EXchanges (DEXs) do not rely on a central authority to hold users’ funds or execute trades. Instead, they use smart contracts running over a public blockchain such as Ethereum, to automate the trading process. DEXs have become increasingly popular reaching a daily trading volume of billions of dollars DEX Tracker - Decentralized Exchanges Trading Volume (2023).
There are two main types of DEXs, namely, order book DEXs and Automated Market Makers (AMMs). Order book DEXs were introduced first but nowadays AMMs are the most common type of DEX. Order book DEXs are similar to traditional order book exchanges. Users place orders to buy or sell assets, and these orders are matched with each other to execute trades. AMMs work by creating liquidity pools, which are essentially large baskets of assets that are held by the exchange. When a user wants to trade an asset, they do not actually trade with another user. Instead, they trade with the liquidity pool. The AMM will then use a predetermined mathematical formula to determine the price of the trade. One of the most widely used AMMs is the Constant Product MM (CPMM). CPMMs use a constant product formula, that ensures that the product of the reserves of two assets in the liquidity pool remains constant. The CPMM’s formula inputs the volume of the trade and the available liquidity and outputs the price of the trade.
A key feature of the CPMM algorithm is that the price it computes reflects the relative scarcity of an asset in the AMM’s liquidity pool. This effectively translates traders’ demand into prices: when market sentiment about an asset’s price rises, traders buy the asset from the AMM, reducing its availability in the liquidity pool until the AMM’s price aligns with market sentiment.
Market fragmentation across multiple AMMs presents challenges, the most notable being the reliance on costly arbitrage to equalize prices. Each CPMM independently adjusts swap rates based on its own trading history, discovering asset prices at different paces. If two CPMMs start with the same exchange rate and one executes a trade, their rates will diverge, with larger trades causing greater price differences. These discrepancies create arbitrage opportunities, systematically exploited by arbitrageurs to restore price consistency. A recent study McLaughlin et al. (2023) estimates that arbitrage profits on the Ethereum blockchain alone amount to $170$ million per year, leading to higher costs for traders and reduced profits for liquidity providers, and thus reducing the efficiency of the blockchain ecosystem.
Another consequence of market fragmentation is increased slippage, i.e. the price’s sensitivity to order size. This effect is particularly pronounced in AMMs with small liquidity pools, where large orders significantly alter asset scarcity, causing sharp price impacts in the CPMM. Clearly, aggregating all liquidity pools into a single AMM would substantially reduce slippage if the CPMM algorithm remains in use.
Standard microeconomic analysis suggests that slippage reduces market efficiency as it deviates from the price-taking assumption; see Aldridge (2022). In practice, however, a more pressing issue is that higher slippage increases the profitability of sandwich attacks. In this front-running strategy, an attacker places a buy order before and a sell order after a victim’s trade, manipulating the price to profit from the slippage they induce. As a result, the victim receives a worse price, while the attacker exploits the price difference for profit.
In this paper, we explore how to address these inefficiencies by computing prices in the spirit of the CPMM but based on the relative scarcity of assets in the aggregate liquidity pools of all AMMs rather than in individual AMM pools. This approach allows prices in a fragmented AMM ecosystem to behave as if they were set by a single AMM.
Our first observation is the failure of the naive approach of applying the CPMM formula using aggregate liquidity pools instead of individual ones, what we call the naive Global Market Maker (nGMM). This algorithm can be easily manipulated by strategically rearranging transactions, potentially leading to the depletion of the AMM reserves.
Next, we characterize algorithms that fall between the standard CPMM and the nGMM but are not manipulable as described in the previous paragraph. Among these, we define the Global Market Maker (GMM) as the algorithm closest to the nGMM.
In the rest of the paper, we study the properties of the GMM and compare it to the CPMM. We show that the GMM eliminates all arbitrage opportunities and that the profits arbitrageurs would earn when all AMMs use the CPMM algorithm instead are shared between the AMMs and the traders.
We also investigate, both theoretically and numerically, how much the GMM reduces the profitability of sandwich attacks compared to the CPMM. In all simulated examples using real data, we observe reductions of over 50%.
Finally, we analyze the performance of the GMM with respect to impermanent loss. This is defined, see Hafner & Dietl (2024), as the loss incurred by a liquidity provider when depositing assets into an AMM, if the relative prices of those assets change. Since the GMM incorporates liquidity information from other AMMs, the price it offers adjusts before reaching the AMM which minimises impermanent losses.
The rest of the paper is structured as follows. Section 2 discuss the related literature. Section 3 introduces the currently most used algorithm, the CPMM, and discuss its shortcomings. Section 4 analyses our propososal of improvement of the design of the CPMM, the GMM, and Section 5 illustrates how the GMM increases performance relative to the GMM. Section 7 concludes.
2 Literature Review
Several alternative mechanisms have been proposed to mitigate MEV and the costs of arbitrage. Canidio & Fritsch (2024) introduce batch trading, a system where orders are executed in batches rather than continuously, thereby reducing arbitrage and sandwich attacks. This mechanism leverages arbitrageurs’ competition to eliminate price manipulation but relies on oracle pricing. Xavier Ferreira & Parkes (2023) study verifiable sequencing rules to prevent MEV and find that while some non-zero miner profits remain, their proposed sequencing rule ensures that user transactions remain unaffected. Canidio & Fritsch (2024) develop a game-theoretic model to differentiate front-running from legitimate trades. Their proposed protocol reduces front-running risk but requires additional message exchanges, which could impact efficiency. Zhou et al. (2021) proposes a modification of AMM algorithms that incorporate optimal on-chain swap routing and arbitrage.
The literature also provides a detailed analysis of the consequences of arbitrage between AMMs. Milionis et al. (2024) identifies arbitrage as a rebalancing cost for LPs in a theoretical model of liquidity provision. Milionis et al. (2023) studies the effect of fees on arbitrage. Mazor & Rottenstreich (2023) explores cross-chain arbitrage in decentralized exchanges, providing an empirical analysis beyond the Ethereum ecosystem. Hansson (2022) examine arbitrage opportunities in Ethereum, identifying \$30 million in cross-exchange and triangular arbitrage profits across 63,168 trading pairs between July 2020 and February 2022.
The profitability of sandwich attacks and more generally the issue of Maximal Extractable Value (MEV) has been widely documented. Daian et al. (2019) introduce the concept of MEV.Canidio & Danos (2024) provide a game-theoretic analysis of sandwich attacks. Qin et al. (2021) estimate that total extractable value over a 32-month period reached \$541 million, highlighting the scale of MEV extraction. Chi et al. (2024) develop methodologies to identify sandwich attacks, finding that \$675 million was extracted before September 2022 and noting that high-volatility tokens are primary MEV targets. Gramlich (2020) provide a comprehensive literature review on MEV, categorizing extraction strategies and countermeasures. Weintraub et al. (2022) analyze MEV extraction in private pools, quantifying the role of flashbots and other sophisticated arbitrage techniques.
More generally, our work is related to the literature on Automated Market Makers (AMMs). Lehar & Parlour (2024) provide a general characterization of AMMs, comparing them to Centralized Exchanges (CEXs). Aoyagi (2023) analyze strategic liquidity provision in AMMs. Cong et al. (2021) analyse the conditions under which CPMM pricing is an optimal liquidity provision mechanism.
3 Our Benchmark: The Constant Product Market Maker (CPMM)
An automated market maker (AMM) is a smart contract that implements a decentralized exchange protocol based on an algorithm. These protocols allow users, known as liquidity providers (LPs), to deposit pairs of assets into liquidity pools, which are then used to execute trades automatically according to the AMM algorithm.
In this section, we describe the algorithm that we use as the starting point of our analysis. This is the most popular algorithm used by AMMs: the constant product market maker (CPMM) introduced by Buterin (2016). Hayden Adams provided the first implementation of the CPMM in his decentralized exchange, Uniswap, in 2017, and Uniswap quickly became the most popular decentralized exchange on Ethereum, with CPMMs emerging as the dominant type of AMM.
The Constant Product Market Maker (CPMM) manages a pool of reserves consisting of two assets. A trader who wishes to swap one asset for another sends a certain amount of the first asset into the AMM. The CPMM algorithm then calculates the amount of the second asset to be returned to the trader, ensuring that the product of the reserves of both assets in the liquidity pool remains constant.
Formally, let $x_{i}$ and $y_{i}$ be the reserves of assets $X$ and $Y$ in the algorithm’s liquidity pool. If a trader sends an amount $\Delta x_{i}>0$ of asset [^1] Since $X$ and $Y$ are arbitrary assets, it is sufficient to describe orders in which the trader sends asset $X$. The case in which the trader sends asset $Y$ can be handled by relabeling the assets. $X$ into the algorithm, the CPMM returns an amount $\Delta y_{i}>0$ of asset $Y$ such that:
$x_{i}\cdot y_{i}=(x_{i}+\Delta x_{i})\cdot(y_{i}-\Delta y_{i}).$ (1)
Solving this equation for $\Delta y_{i}$ allows for a characterization in terms of the amount of $Y$ obtained by a trader that sends $\Delta x_{i}$ units of $X$ to the CPMM:
$\Delta y_{CPMM}(\Delta x_{i};x_{i},y_{i})=\frac{y_{i}}{x_{i}+\Delta x_{i}}\Delta x_{i}=\frac{r_{i}}{1+\frac{\Delta x_{i}}{x_{i}}},$ (2)
where $r_{i}\equiv\frac{y_{i}}{x_{i}}$ is the ratio of reserves of the AMM. The interpretation of the last equation is that the terms of trade $\frac{\Delta y}{\Delta x}$ of the CPMM are given by a measure of the relative scarcity of the assets in the liquidity pool, adjusted by a measure of the order’s size relative to the liquidity pool of the AMM. The first term represents the marginal price, as it corresponds to the limit of the terms of trade when $\Delta x_{i}$ tends to zero. The second term captures slippage, which quantifies the rate at which the terms of trade deteriorate as the order size increases relative to the liquidity pool.
We identify next some drawbacks of the CPMM when it operates in a fragmented ecosystem with several AMMs, all using the CPMM algorithm. The first and most obvious is that differences in prices arise naturally if traders do not split the orders optimally between the different AMMs. This creates profitable arbitrage opportunities that become a cost for traders as we illustrate in the next example.
{exampleboxinner}
[frametitle=Toy example - Part 1: Arbitrage between CPMMs.] Consider two AMMs that use the CPMM algorithm, each with a liquidity pool containing 100 ETH and 400,000 UST. The marginal price is the same for both CPMMs and equal to 4,000 UST/ETH. All the computations in this example are done using (2).
Suppose that a trader requires 10ETH. He can get them from AMM 1 sending 44,444 UST to AMM 1. After the trade, AMM 1’s pool will contain 90 ETH and 444,444 UST. The marginal price in AMM 1 is now 4,938 UST/ETH. An arbitrage opportunity has emerged.
An arbitrageur can profit by sending 5 ETH to AMM 1 to get 23392 UST and then sending 21,052 UST to AMM 2 to obtain 5 ETH back. After the arbitrage, the reserves of both CPMMs will be equal to 95 ETH plus 421,052 UST and the arbitrageur makes a profit of 2,339 UST.
Suppose that after this trade, the trader wants to convert the 10 ETH back into UST. He can send them to AMM 1 to obtain 40,100 UST. Thus the AMM 1 pool will contain 105 ETH and 380,952 UST. Again, an arbitrage operation has emerged which is seized by an arbitrageur who sends 5 ETH to AMM 2 and then sends the UST received to AMM 2. These two trades give the arbitrageur a profit of 2,005 UST. After these two arbitrage operations, the reserves of both AMMs will be restored to 100 ETH and 400,000 UST each.
Overall, in these 6 operations, the CPMMs retained their initial reserves intact, the arbitrageur has made a profit of 4,344 UST which matches to the net cost to the trader.
Thus, whereas arbitrage provides the service of equalising the marginal prices across AMMs, it is at a cost for the trader(s) which is equal to the profits of the arbitrageurs.
The second implication of the fragmentation of the reserves into different AMMs is that each individual AMM only has a portion of total liquidity reserves which makes slippage substantially larger than in a hypothetical AMM holding the aggregate reserves of all AMMs. Higher slippage makes it more profitable to conduct a sandwich attack. This is a manipulative attack that combines two trades, one before the target transaction (frontrunning) and another one after (backrunning) to profit from the price movement caused by the victim’s trade. In CPMMs, this involves inserting trades around a large swap to capitalize on slippage, forcing the victim to execute their trade at a worse price while the attacker profits from the price difference. We illustrate sandwich attacks in the next example.
{exampleboxinner}
[frametitle=Toy example - Part 2: MEV extraction in CPMM.] Consider an AMM using the CPMM algoritm and with a liquidity pool containing 100 ETH and 400,000 UST. The marginal price of ETH is then 4,000 UST/ETH. Suppose that a trader interested in converting 40,000 UST in ETH sends them to the AMM. The MEV-extractor can perform a sandwich attack by sending 60,000 UST just before the victim’s transaction is executed to send back immediately after the ETH obtained. This would work as follows:
  • Frontrunning: MEV-extractor transaction 1 sends 60,000 UST to get 13.0435 ETH in return, and changes the vector of reserves of AMM 1 to 86,9565 ETH and 460.000 UST. The marginal price of ETH increases to 5,290 UST/ETH.
  • Victim’s transaction: The victim sends 40,000 UST to obtain in return 6.9565 ETH. The resulting reserves of the AMM are then 80 ETH and 500,000 UST. Note that in the absence of the frontrunning transaction, the victim would have received 9,0909 ETH, or equivalently, they only needs to send 29,906 UST to get 6,9565 ETH, i.e. they pay an overpriced of 10,094 UST.
  • Backrunning: MEV-extractor transaction 2 sends 13.0435 ETH to get 70.094 UST. The AMM’s reserves after the trade are 110 ETH and 363,636 UST. The MEV-extractor obtains a total profit of 10,094 UST, which is equal to the trader’s overprice.
A simple analytical exercise allow us to quantify the profits derived by a general MEV sandwich operation consisting of submitting a frontrunning order $\hat{\Delta}x$ to a traders order $\Delta x$ followed by a backrunning order in which the manipulator sends back to the AMM the amount obtained in the frontrunning trade. We refer to this manipulative operation as an elementary MEV sandwich $\hat{\Delta}x$ to order $\Delta x$. The next result provides the corresponding analytical formula for the profits.
Proposition 1.
The profit of an elementary MEV sandwich $\hat{\Delta}x$ to a trader order $\Delta x$ to an AMM i with liquidity reserves $(x_{i},y_{i})$ and running the CPMM algorithm is equal to:
$\left(\frac{\left(1+\frac{\hat{\Delta}x}{x_{i}}+\frac{\Delta x}{x_{i}}\right)^{2}}{\left(1+\frac{\hat{\Delta}x}{x_{i}}+\frac{\Delta x}{x_{i}}\right)\left(1+\frac{\hat{\Delta}x}{x_{i}}\right)-\frac{\Delta x}{x_{i}}}-1\right)\hat{\Delta}x$ (3)
Proof.
Since the proof of this proposition is a particular case of Proposition for the case $x_{i}=x$ we postpone the proof. ∎
Figure 1 illustrates the proposition for different values of $\hat{\Delta}x$.
Figure 1: The graph plots the profits of elementary MEV sandwich manipulations $\hat{\Delta}x$ for $x_{i}=400,000$ and $\Delta x=40,000$.
Finally, the fragmentation of the AMM ecosystem, together with the CPMM algorithm, means that AMMs often operate with outdated prices, leading to larger impermanent losses. This is the loss liquidity providers may experience when supplying liquidity to AMMs due to the automatic rebalancing of token holdings in the liquidity pool as the relative price of tokens fluctuates. Specifically, as one token’s price decreases relative to the other, liquidity providers end up holding more of the depreciated asset and less of the appreciated one than they initially deposited. The term impermanent reflects that the loss may be reversed if the price ratio returns to its original state. However, in practice, such reversals are rare, making impermanent loss a crucial factor for liquidity providers to consider when engaging with AMMs.
Mathematically, impermanent loss is defined as the relative loss incurred by an LP compared to simply holding the original asset allocation: [^2] See Lipton et al. (2024) for a comprehensive discussion of alternative measures of impermanent losses. Ours correspond to what they call relative IL of borrowed Liquidity Provision.
$\text{IL}_{i}=1-\frac{V_{i}}{V^{0}_{i}}$
where $V_{i,t}\equiv y_{i}+rx_{i}$ is the value of the liquidity holdings of the AMM at the current price, and $V^{0}_{i}\equiv y_{i}^{0}+rx_{i}^{0}$ is the value of the liquidity holdings if the liquidity provider had kept the initial asset allocation. The impermanent losses for the CPMM has a simple expression under the assumption that the initial ratio of reserves of the AMM and the final ratio of reserves are equal to $r_{0}$ and $r$, respectively, and so are the initial and final marginal prices of the AMM:
$\text{IL}_{i}=1-\frac{2}{\sqrt{\frac{r}{r_{0}}}+\sqrt{\frac{r_{0}}{r}}}$
where $r_{0}$ and $r$ are the initial and final price ratios of the token pair, respectively. Figure 2 provides a graph of the impermament losses as a function of the ratio $\frac{r}{r_{0}}$ and the next example illustrates the concept.
Figure 2: The graph plots the impermanent losses as a function of the price increase $\frac{r}{r_{o}}$.
{exampleboxinner}
[frametitle=Toy example - Part 3: Impermanent loss in CPMM.] Consider two automated market makers AMM 1 and AMM 2 using CPMM algorithm, each with a liquidity pool with 100 ETH and 400000 UST. The marginal price is then 4,000 UST/ETH.
Suppose that an event makes traders in the market believe that the new price should be 3,000 UST/ETH. As a result, they sell ETH to each of the two AMMs until the marginal price moves to 3,000 UST/ETH in both CPMMs. The new reserves for each CPMM after the transactions are $115.47$ ETH and $346,410.16$ UST.
We compute the impermanent loss of each AMM as the difference between the value of the CPMM’s original reserves valued at the new price ($V_{i}^{0}$) and the current reserves valued at the new price ($V_{i}$) normalised by the former value, i.e.,
[…]
$IL_{i}$ $=$ $\frac{V_{i}^{0}-V_{i}}{V_{i}^{0}}=\frac{700,000UST-692,820.16UST}{700,000UST}=1.03\%.$
4 The Analysis of Global Market Makers
In this section, we study how to twist the design of the CPMM to solve or minimize the problems we discussed in the previous section. Our proposal will be based on a very simple observation. Each AMM has available in blockchains like Ethereum information about the liquidity pools of all other AMMs and this information could be used to determine prices as if the market were not fragmented. We refer these algorithms as global algorithms and we shall focus on global algorithms with certain properties and study their performance.
[…]
$x\equiv\sum_{j\in I}x_{j},\quad\text{and}\quad x_{-i}\equiv\sum_{j\in I\setminus\{i\}}x_{j},$
with analogous definitions for $y$ and $y_{-i}$. Let also $r=\frac{y}{x}$ be the global ratio of reserves.
We are interested in algorithms that determine the terms of trade usign the information contained in the vector of liquidity reserves or all AMMs. We refer to such algorithms as global. The global algorithm that mimics the behavior of a CPMM algorithm with the aggregate reserves is as follows.
Definition: Given a vector of liquidity pools $(x,y)$, the naive Global Market Maker (nGMM) algorithm swaps $\Delta x_{i}>0$ of $X$ for an amount of $Y$:
$\Delta y_{\text{nGMM}}(\Delta x_{i};x,y)=\frac{y}{x+\Delta x_{i}}\Delta x_{i}=\frac{r}{1+\frac{\Delta x_{i}}{x}}\Delta x_{i},$ (4)
if less than $y_{i}$, and $\Delta y_{\text{nGMM}}(\Delta x_{i};x,y)=y_{i}$, otherwise.
This algorithm determines the terms of trade as if executing a hypothetical CPMM with reserves equal to the sum of individual reserves. Consequently, the terms of trade reflect both global scarcity and global slippage. Notably, if all AMMs adopt this algorithm, arbitrage opportunities are eliminated, as illustrated in the following example, and transformed into better terms of trade for traders.
{exampleboxinner}
[frametitle=Toy example - Part 4: nGMM] Consider two automated market makers AMM 1 and AMM 2 using CPMM algorithm, AMM 1’s pool contains 90 ETH and 444,444 UST and AMM 2’s pool 100 ETH and 400000 UST. This is exactly the distribution of reserves in Part 1 before arbitrage is executed.
However, no strictly profitable arbitrage opportunities exist here: if an arbitrageur sends 5 ETH to AMM 2 gets 21,652 UST that after sending them to AMM 1 delivers 5 ETH. Thus, zero net trade.
Suppose instead that the a trader sends 10 ETH to AMM 1 when AMM 1’s pool contains 90 ETH and 444,444 UST and AMM 2’s pool 100 ETH and 400000 UST. The trader gets 42,222 UST in return. This is 2,122 ETH more than in the CPMM of Part 1 after arbitrage.
However, the nGMM has two significant drawbacks. The first drawback is immediate: since nGMM sets prices based on the aggregate liquidity, it can lead to complete depletion of one asset’s reserves, forcing the AMM to cease operations. This issue does not arise in the CPMM by design: withdrawing an amount $\Delta y$ of $Y$ arbitrarily close to $y_{i}$ requires supplying an unbounded amount of $X$, causing the per-unit price of $Y$ to diverge to infinity.
The second drawback is more subtle: the nGMM can be exploited if some particular sequences of transactions appear in the market. Specifically, the algorithm may sell an asset at a low price when it is relatively abundant across other AMMs, only to repurchase it at a higher price when it becomes scarce in later trades. As a result, the AMM may suffer a net decrease in its holdings of both assets.
We illustrate these two vulnerabilities with an example.
{exampleboxinner}
[frametitle=Toy example - Part 5: Exhausting and Exploiting nGMM] Consider two automated market makers AMM 1 and AMM 2 using nGMM, each with a liquidity pool with 100 ETH and 400000 UST.
  • Exhausting nGMM. Suppose that a trader sends 800,000 UST to AMM 1. According to the nGMM formula, AMM 1 returns 100 ETH to the trader, fully depleting its ETH reserves.
  • Exploiting nGMM. We start again with the initial distribution of reserves of 100 ETH and 400,000 UST in both AMM 1 and AMM 2. Consider the effect of the following transactions:
    • Transaction 1: Trade A sends 10 ETH to AMM 1 and gets 38,095 UST in return. The resulting reserves of AMM 1 are 110 ETH and 361,905 UST. AMM 2 still has 100 ETH and 400,000 UST.
    • Transaction 2: Trade B sends 10 ETH to AMM 2 and gets 34,632 UST in return. The resulting reserves of AMM 2 are then 110 ETH and 365,368 UST. AMM 1 are 110 ETH and 361,905 UST.
    • Transaction 3: Trade C sends 38,095 UST to AMM 1 to get 10.95 ETH. The reserves for AMM 1 would then be 99,05 ETH and 400,000 UST, which is strictly less than its initial reserves (same amount of UST and less of ETH). AMM 2 remains with reserves 110 ETH and 365,368 UST.
    • Transaction 4: Trade D sends 34,632 UST to AMM 2 to get 9.05 ETH in return. The reserves for AMM 2 would then be 100.95 ETH and 400,000 UST, which is strictly greater than its initial reserves (same amount of UST and greater of ETH).
    AMM 2 gains 0.95 ETH that have been obtained from AMM 1. The net variation of trades A, B, C and D is equal to zero. Thus, AMM 2 can still 0.95 ETH from AMM 1 submitting the trades A, B, C and D.
Consequently, the nGMM is not only vulnerable to exploitation, but other AMMs have incentives to do so by submitting zero net-cost orders.
These two vulnerabilities render the nGMM unsuitable for practical implementation. Our approach, therefore, is to explore how much we can adjust the CPMM to approximate the nGMM while avoiding these vulnerabilities.
For this purpose, we adopt a highly conservative approach. We restrict our attention to algorithms that, once implemented by an AMM, say $i$, prevent exhaustion and exploitation for any initial distribution of reserves $(\vec{x}^{0},\vec{y}^{0})$ and any sequence of swaps.
Formally, a sequence of swaps is given by $\{\iota(t),\Delta x_{t},\Delta y_{t}\}_{t=1}^{T}$, where the index $t$ represents the order of transactions, $\iota(t)$ denotes the identity of the AMM handling the swap, and $\Delta x_{t}$ and $\Delta y_{t}$ represent the quantity of $X$ sent to and the quantity of $Y$ returned by AMM $\iota(t)$, respectively. The case in which the AMM receives $Y$ and supplies $X$ is handled by allowing $\Delta x_{t}$ and $\Delta y_{t}$ to take negative values.
We impose only minimal consistency requirements. We say that a sequence of transactions $\{(\iota(t),\Delta x_{t},\Delta y_{t})\}_{t=1}^{T}$, where $\Delta x_{t}>0$, is consistent with the global algorithm of AMM $i$ if at any $t$ such that $i=\iota(t)$, swapping $\Delta x_{t}$ for $\Delta y_{t}$ (or vice versa) is compatible with the algorithm of AMM $i$ and the vector of liquidity reserves updated with the preceding swaps $\{(\iota(s),\Delta x_{s},\Delta y_{s})\}_{s=1}^{t-1}$.
Definition: The global algorithm used by AMM $i\in I$ is non-exhaustible if there exists no distribution of liquidity pools and no sequence of swaps consistent with the algorithm of $i$ that causes AMM $i$ to run out of either asset, i.e., such that $x_{i}=0$ or $y_{i}=0$.
Definition: A global algorithm employed by AMM $i$ is exploitable if there exists an initial distribution of liquidity pools $(x^{0}_{i},y^{0}_{i})$ and a sequence of swaps consistent with AMM $i$’s algorithm such that AMM $i$ ends up with a lower liquidity pool, i.e., $(x_{i},y_{i})\leq(x_{i}^{0},y_{i}^{0})$. [^3] We denote $(x_{i},y_{i})\leq(x_{i}^{\prime},y_{i}^{\prime})$ if $x_{i}\leq x_{i}^{\prime}$ and $y_{i}\leq y_{i}^{\prime}$, with at least one inequality being strict.
[…]
The restriction to algorithms between the CPMM and the nGMM should be interpreted as a minimal departure from existing practices. The CPMM represents the current state of the art, while the nGMM serves as its ideal generalization to mitigate the inefficiencies caused by market fragmentation. Our objective is to approximate the nGMM as closely as possible while deviating as little as possible from the CPMM.
Proposition 2.
A global algorithm in between the CPMM and the nGMM is not exploitable if and only if it satisfies:
$\Delta y(\Delta x_{i};\vec{x},\vec{y})\leq\Delta y_{CPMM}(\Delta x_{i};x_{i},y_{i}),$ (5)
[…]
Proof.
By contradiction, suppose a distribution of reserves $(\vec{x}^{0},\vec{y}^{0})$ and a swap that sends $\Delta x_{i}>0$ of $X$ and gets $\Delta y_{i}$ units of $Y$ in return and that violates (5). The definition of CPMM means that the swap $\Delta x_{i}$ decreases the product of reserves, this is,
$x_{i}^{0}y_{i}^{0}>x_{i}y_{i},$ (6)
for $x_{i}\equiv x_{i}^{0}+\Delta x_{i}$ and $y_{i}\equiv y_{i}^{0}-\Delta y_{i}$. Now suppose that there is a sequence of swaps submitted to the other AMMs that leads to a new vector of reserves $(x_{-i},y_{-i})$ such that:
$\frac{y_{-i}}{x_{-i}}\leq\frac{y_{i}}{x_{i}},$ (7)
and consider a swap that sends $\Delta y_{i}$ of $Y$ to AMM $i$. Then:
$\Delta x_{CPMM}(\Delta y_{i};x_{i},y_{i})$ $=$ $\frac{\frac{x_{i}}{y_{i}}}{1+\frac{\Delta y_{i}}{y_{i}}}\Delta y_{i}$
$\leq$ $\frac{\frac{x}{y}}{1+\frac{\Delta y_{i}}{y}}\Delta y_{i}$
$=$ $\Delta x_{nGMM}(\Delta y_{i};x,y),$
where the two equalitites follow from (2) and (4), and the inequality from the fact that $x_{i}<x$ and that (7) implies that $\frac{y_{i}}{x_{i}}\geq\frac{y}{x}$. This together with our restriction to global algorithms between the CPMM and the nGMM means that the amount of $X$ that AMM $i$ gives to the trader is at least:
$\hat{\Delta}x_{i}\geq\Delta x_{CPMM}(\Delta x_{i};x_{i},y_{i}),$ (8)
which by the definition of $\Delta x_{CPMM}(\Delta x_{i};x_{i},y_{i})$ implies that:
$(x_{i}-\hat{\Delta}x_{i})(y_{i}+\Delta y)\leq x_{i}y_{i}.$
Since $y_{i}=y^{0}_{i}-\Delta y_{i}$, (6) means that $x_{i}-\hat{\Delta}x_{i}<x_{i}^{0}$ and thus the updated vector of reserves of AMM $i$ satisfies,
$(x_{i}-\hat{\Delta}x_{i},y^{0}_{i})\leq(x_{i}^{0},y^{0}_{i}),$
as desired. ∎
Thus, global algorithms in between the CPMM and nGMM buy asset $X$ at a price no larger than the CPMM price to avoid exploitation. Intuitively, the CPMM price guaranties that when the AMM buys an asset and sells it back the vector of reserves returns to its initial level. Thus, whenever a global algorithm buys one of the assets more expensive than the CPMM it is going to be exploited if it sells it back at CPMM prices or lower. But this will always happen with global algorithms with prices in between the CPMM and nGMM if the asset becomes so abundant among the other AMMs between the initial sale and the sell back that the nGMM price becomes lower than the CPMM price.
Furthermore, since the CPMM is not exhaustible and Proposition 2 says that to avoid exploitation the price at which the algorithm buys must be at most the CPMM price, we can conclude that the algorithms that satisfy Proposition 2 are not exhaustible.
Corollary 1.
A non exploitable global algorithm between the CPMM and the nGMM is also non exhaustible.
Since CPMM prices put a lower bound to global non exploitable algorithms, the closer one can get to nGMM prices while avoiding exploitation (and thus exhaustability) is to use nGMM prices unless bound by the CPMM.
Definition: The Global Market Maker (GMM) algorithm is defined by:
$\Delta y_{GMM}(\Delta x;x,y)=\min\{\Delta y_{CPMM}(\Delta x;x_{1},y_{1}),\Delta y_{nGMM}(\Delta x;x,y)\}$ (9)
To understand the GMM, it is useful to distinguish whether a swap $\Delta x_{i}>0$ moves the AMM’s reserve ratio away from that of the other AMMs, which we call a divergent swap. This occurs when the initial vector of reserves satisfies $\frac{y_{i}}{x_{i}}\leq\frac{y_{-i}}{x_{-i}}$. In this case, asset $X$ is relatively less scarce locally than globally, causing the nGMM to pay a higher price $\Delta y^{nGMM}_{i}$ for $X$ than the CPMM price $\Delta y^{CPMM}_{i}$. Consequently, the GMM price $\Delta y^{GMM}_{i}$ coincides with the CPMM price. This is illustrated in Figure 3.
Figure 3: The graph illustrates a divergent swap where the GMM receives $\Delta x_{i}$ units of $X$ and returns $\Delta y_{i}^{GMM}$ of $Y$ to the trader, compared to the returns in the CPMM ($\Delta y_{i}^{CPMM}$) and nGMM ($\Delta y_{i}^{nGMM}$) cases. It also shows the possible liquidity reserves of AMM $i$ after a swap, depending on the algorithm used (CPMM, nGMM, GMMM).
A swap $\Delta x_{i}>0$ is non-divergent when $\frac{y_{i}}{x_{i}}>\frac{y_{-i}}{x_{-i}}$. There are two possible cases. The first, which we refer to (with a slight abuse of terminology) as a convergent swap, occurs when the nGMM returns less $Y$ than the CPMM, causing the GMM to coincide with the nGMM. This happens when $\Delta x_{i}$ is small enough that the swap moves the reserve ratio toward that of the other AMMs, justifying the name. However, it can also occur if $\Delta x_{i}$ is large enough to decrease the reserve ratio below that of the other AMMs but not so large that the nGMM price exceeds the CPMM price. This is illustrated in Figure 4.
Figure 4: The graph illustrates a convergent swap where the GMM receives $\Delta x_{i}$ units of $X$ and returns $\Delta y_{i}^{GMM}$ of $Y$ to the trader, compared to the returns in the CPMM ($\Delta y_{i}^{CPMM}$) and nGMM ($\Delta y_{i}^{nGMM}$) cases. It also shows the possible liquidity reserves of AMM $i$ after a swap, depending on the algorithm used (CPMM, nGMM, GMMM).
The last case, the overshooting swap, is when the non divergent swap $\Delta x$ is so large that the nGMM pays a higher price $\Delta y^{nGMM}_{i}$ for $X$ than the CPMM price $\Delta y^{CPMM}_{i}$, causing the GMM price $\Delta y^{GMM}_{i}$ to match the CPMM price. [^4] Formally, the overshooting case occurs at the point where: $\frac{y_{i}}{x_{i}}\cdot\frac{y_{i}-\Delta y_{CPMM}(\Delta x_{i};x_{i},y_{i})}{x_{i}+\Delta x_{i}}=\frac{y}{x}\cdot\frac{y-\Delta y_{CPMM}(\Delta x_{i};x_{i},y_{i})}{x+\Delta x_{i}}.$ We refer to this case as an overshooting swap and illustrate it in Figure 5.
Figure 5: The graph illustrates an overshooting swap where the GMM receives $\Delta x_{i}$ units of $X$ and returns $\Delta y_{i}^{GMM}$ of $Y$ to the trader, compared to the returns in the CPMM ($\Delta y_{i}^{CPMM}$) and nGMM ($\Delta y_{i}^{nGMM}$) cases. It also shows the possible liquidity reserves of AMM $i$ after a swap, depending on the algorithm used (CPMM, nGMM, GMMM).
Finally, note that the GMM always receives prices that are not worse than the CPMM, strictly better in the case of convergent swaps. This means that the CPMM product of reserves is expected to grow since the CPMM algorithm keeps the product constant. Next corollary formalises this idea.
Corollary 2.
The product of reserves of an AMM with the GMM algorithm weakly increases with every swap, strictly if the swap is convergent.
5 Properties of the GMM
In this section, we show that by exploiting efficiently the aggregation of the information from the liquidity pools of the other AMMs, the GMM improves in the following key dimensions of the design of an AMM: first, it eliminates arbitrage opportunities and splits the efficiency gains between traders and AMMs; second, it reduces slippage and thus the profitability of MEV sandwich trades, and third, it reduces impermanent losses. A second order effect is that the share of arbitrage profits that end in the AMMs increase their liquidity pools and as a consequence further reduce their slippage.
5.1 Arbitrage
We revisit Example 1 to show that in the GMM no arbitrage opportunities appear, and that the profit that arbitrageurs obtain in the CPMM is split between the AMMs and the traders.
{exampleboxinner}
[frametitle=Toy example - Part 6: GMM]
Consider two AMMs that use the GMM algorithm, each with a liquidity pool containing 100 ETH and 400,000 UST.
A trader that sends 44,444 UST to AMM 1 gets 10 ETH in return, which is the minimum between the CPMM return of 10 ETH, and the nGMM return of 10.53 ETH. This is also consistent with the fact that since both GMMs have the same reserves, the trade is divergent. After this trade, as in part 1, GMM1’s pool updates to 90 ETH and 444,444 UST and GMM2’s pool remains unchanged.
No arbitrage opportunities exist because arbitrage requires convergent trades which are priced using the nGMM algorithm. For instance, an arbitrageur that sends 5 ETH to AMM 1, gets the minimum between the CPMM return of 23,392 UST and the nGMM return of 21652 UST. This trade updates AMM 1 pool of reserves to 95 ETH and 422,792 UST. If after that the arbitrageur sends the 21652 UST obtained to AMM 2 gets the minimum of the CPMM return of 5.13 ETH and the nGMM return of 5 ETH. Thus, the arbitrageur’s net trade is zero, and no profit is made.
Suppose now that, similarly to Example 1, the reverse transaction is executed, namely that a trader sends 10 ETH to AMM 1. The trade is convergent and the AMM returns to the trader 42,222 UST and the resulting reserves of GMM2 after the trade are 100 ETH and 402,222 UST.
After this trade, the reserves of the GMM have increased in 2,222 UST and the trader paid 2,222 UST less than in the CPMM with arbitrage described in Part 1. In other words, GMM splits evenly the profit of the arbitrageur between the trader and the GMM.
The insights of the above example hold true in general.
Proposition 3.
There are no striclty profitable arbitrage opportunities between a set of AMMs that use the GMM.
Proof.
Arbitrage consists of a sequence of swaps initiated by the arbitrageur across different AMMs. Regardless of which AMM executes each swap, the GMM always offers worse terms of trade to the arbitrageur than the nGMM. Moreover, the nGMM preserves the product of aggregate reserves. Consequently, we have:
$xy$$\leq$ $(x+\Delta x_{1})(y-\Delta y_{1})$
$\leq$ $(x+\Delta x_{1}+\Delta x_{2})(y-\Delta y_{1}-\Delta y_{2})$
$\leq$ $\dots$
$\leq$ $\left(x+\sum_{t=1}^{T}\Delta x_{t}\right)\left(y-\sum_{t=1}^{T}\Delta y_{t}\right),$
for any sequence of trades sent by the arbitrageur to the different AMMs, where $\Delta x_{i}$ represents the amount of $X$ moved from the arbitrageur to an AMM, and $\Delta y_{i}$ represents the amount of $Y$ received by the arbitrageur from the AMM.
This inequality implies that either $\sum_{t=1}^{T}\Delta y_{t}\leq 0$ or $\sum_{t=1}^{T}\Delta x_{t}\geq 0$, meaning that the arbitrageur’s net position after the swaps,
$\left(-\sum_{t=1}^{T}\Delta x_{t},\sum_{t=1}^{T}\Delta y_{t}\right),$
must be either zero or have a strictly negative component. Thus, there are no strictly profitable arbitrage opportunities. ∎
Thus, replacing the CPMM with the GMM could benefit both traders and AMMs by converting strictly profitable arbitrage opportunities into additional profits for them. For any sequence of swaps submitted by traders or arbitrageurs to different AMMs, the total assets transferred to and withdrawn from the AMMs must equal the total variation in liquidity reserves. Arbitrage profits create a wedge between the net flows of traders and AMMs. Eliminating arbitrage removes this wedge, allowing for higher profits for both.
Furthermore, if traders strictly benefit from the switch, their terms of trade must improve, whereas if AMMs strictly benefit, their liquidity pools should grow, reducing slippage and improving terms of trade for future traders. In either case, better terms of trade should lead to increased trading and thus higher potential gains from the switch.
To assess the distribution of gains, we consider a stylized benchmark. First, we assume that all strictly profitable arbitrage opportunities are exploited by arbitrageurs and that traders always submit their trades to the AMM offering the most favorable terms of trade, a strategy we call optimal routing. As we shall see, this assumption aligns with empirical evidence.
Second, when all AMMs use the CPMM algorithm, the individual liquidity ratios after maximum arbitrage, $\frac{y_{i}}{x_{i}}$, must be equal across AMMs. However, this common ratio may depend on how arbitrageurs extract their profit. For instance, if arbitrageurs structure their swaps to realize profits exclusively in $Y$, ensuring that the net position of asset $X$ across all arbitrage swaps remains zero, the global liquidity ratio $\frac{\sum_{i}y_{i}}{\sum_{i}x_{i}}$ will decrease, leading to a corresponding decline in the common individual liquidity ratios $\frac{y_{i}}{x_{i}}$.
To abstract from this complication, we assume that arbitrage profit extraction preserves the global liquidity ratio, ensuring that the resulting individual liquidity ratios remain equal to the original global liquidity ratio. We refer to this assumption as balanced arbitrage.
Proposition 4.
Consider an initial distribution of liquidity $(\vec{x},\vec{y})$ and a swap $\Delta x>0$ of $X$ that is optimally routed and compare two scenarios:
  1. (a)
    All AMMs use the CPMM, and all strictly profitable arbitrage opportunities in the initial liquidity distribution are exploited by balanced arbitrage before the swap $\Delta x$ is executed.
  2. (b)
    All AMMs use the GMM algorithm.
(a) provides the trader strictly better terms of trade than (b) if and only if:
$\frac{r}{1+\frac{\Delta x}{x}}>\frac{r_{i}}{1+\frac{\Delta x}{x_{i}}}<\frac{r}{1+\frac{\Delta x}{\sqrt{\frac{\max\{x_{j}y_{j}\}_{j\in I}}{r}}}},\,\forall i\in I.$ (10)
Proof.
We begin with some remarks that apply to both parts of the proof. In (a), the CPMM mechanism and balanced arbitrage imply that AMM $j$’s reserves $(x_{j}^{\prime},y_{j}^{\prime})$ satisfy the equations:
$x_{j}^{\prime}y_{j}^{\prime}$ $=$ $x_{j}y_{j},$
$\frac{x_{j}^{\prime}}{y_{j}^{\prime}}$ $=$ $r,$
which yield the solutions: $x_{j}^{\prime}=\sqrt{\frac{x_{j}y_{j}}{r}}$ and $y_{j}^{\prime}=\sqrt{rx_{j}y_{j}}$. Consequently, the terms of trade for a swap $\Delta x>0$ sent to AMM $j$ are:
$\frac{\sqrt{rx_{j}y_{j}}}{\sqrt{\frac{x_{j}y_{j}}{r}}+\Delta x}=\frac{r}{1+\frac{\Delta x}{\sqrt{\frac{x_{j}y_{j}}{r}}}},$ (11)
where the optimal strategy for the trader is to swap with the AMM $j$ that maximizes $x_{j}y_{j}$. This implies that the last term of (10) represents the most favorable terms of trade for the trader in (a).
In (b), the terms of trade for a swap $\Delta x$ sent to AMM $i$ are:
$\min\left\{\frac{r}{1+\frac{\Delta x}{x}},\frac{r_{i}}{1+\frac{\Delta x}{x_{i}}}\right\}.$ (12)
We now use (11) and (12) to prove the necessary and sufficient conditions, beginning with the ”if” direction. The first inequality in (10) implies that the terms of trade in (b) are determined by the second term in (12). Combining this with the second inequality in (10) and (11) yields the desired result.
For the ”only if” direction, we proceed by contradiction starting with the first inequality in (10). Suppose there exists some $i\in I$ such that:
$\frac{r}{1+\frac{\Delta x}{x}}\leq\frac{r_{i}}{1+\frac{\Delta x}{x_{i}}}.$
This inequality, along with (11) and (12), implies that to complete the contradiction argument, it suffices to show:
$x\geq\sqrt{\frac{x_{j}y_{j}}{r}}\hskip 14.22636pt\forall j\in I,$
To check that this inequality holds true it is replace $r=\frac{y}{x}$ to get:
$\sqrt{xy}\geq\sqrt{x_{j}y_{j}}\hskip 14.22636pt\forall j\in I,$
which is clearly satisfied as desired.
To conclude the proof, suppose now that
$\frac{r}{1+\frac{\Delta x}{x}}>\frac{r_{i}}{1+\frac{\Delta x}{x_{i}}}\geq\frac{r}{1+\frac{\Delta x}{\sqrt{\frac{\max\{x_{j}y_{j}\}_{j\in I}}{r}}}},\,\forall i\in I.$
The first inequality and (12), mean that the most convenient terms of trade for the trader in (b) are:
$\max_{i\in I}\frac{r_{i}}{1+\frac{\Delta x}{x_{i}}},$
which by the second inequality (11) are more beneficial to the trader than the most convenient terms of trade in (a), as desired. ∎
To provide an intuitive interpretation of the Proposition, note that, in both the CPMM and GMM algorithms, the terms of trade depend on the ratio of reserves and slippage. For instance, in the case of the CPMM, when there is no arbitrage or non-convergent GMM orders, a swap $\Delta x$ delivers:
$\frac{r_{i}}{1+\frac{\Delta x}{r_{i}}}\Delta x.$
In the case of convergent orders under the GMM (priced as in the nGMM algorithm), the swap delivers:
$\frac{r}{1+\frac{\Delta x}{r}}\Delta x.$
Similarly, one can show that the CPMM after balanced arbitrage delivers:
$\frac{r}{1+\frac{\Delta x}{\sqrt{\frac{x_{j}y_{j}}{r}}}}\Delta x.$
Thus, the first inequality in (10) requires that all AMMs price GMM orders using CPMM prices, meaning there are no possible convergent swaps for $\Delta x$. If this holds, the second inequality ensures that GMM prices are worse for traders than CPMM prices with arbitrage. This occurs because arbitrage improves the best CPMM price for traders. Clearly, while these conditions can be met in reality, they are very demanding.
The conditions will not hold in the following cases:
  • All AMMs have the same reserve ratio. In this case, there is no room for arbitrage in (a) and no convergent orders in (b), meaning there is no difference between the prices offered by the GMM and the CPMM. [^5] Formally, if $r_{i}=r$, then $\sqrt{\frac{x_{j}y_{j}}{r}}=\sqrt{\frac{x_{j}y_{j}}{r_{j}}}=x_{j}$, which violates the second inequality in (10) for any $i$.
  • All AMMs have different reserve ratios, but $\Delta x$ is small. Heterogeneous reserve ratios and a small order size mean that some AMMs will display a convergent swap in (b). But convergent swaps always offer better prices to traders than CPMM prices with balanced arbitrage: the reserve ratios remain the same, but slippage is lower because the GMM uses as a reference aggregate reserves. [^6] In the limit as $\Delta x\to 0$, (10) fails, as it converges to $r>r_{i}<r$ for all $i\in I$.
  • All AMMs have the same product of reserves, i.e., $x_{j}y_{j}$ is constant across $j$. Again, convergent swaps offer better prices to traders than CPMM prices with balanced arbitrage, so (a) can be better only when (b)´s best prices for traders are CPMM’s prices. But in this case and in both scenarios, AMMs differences can only arise because of differences in reserve ratios. Arbitrage eliminates them in (a), whereas the heterogeneity of reserve ratios in (b) allows traders to find better deals. [^7] A constant product $x_{j}y_{j}$ means that for any $l\in I$, $\sqrt{\frac{\max\{x_{j}y_{j}\}_{j\in I}}{r}}=\sqrt{\frac{x_{l}y_{l}}{r}}=x_{l}\sqrt{\frac{r_{l}}{r}}$, which is strictly less than $x_{l}$ for any $r_{l}<r$. Thus, the second inequality in (10) fails.
For cases where (10) applies and thus worse terms of trade may impact trading activity, there is a natural solution, which we describe next.
5.2 The GMM with Rebalancing
First, an important aspect of the GMM that we have not discussed but is now relevant is how swaps between AMMs using the GMM algorithm are priced. Up to this point, we have implicitly assumed that these trades do not occur. While such orders do not require special consideration under the CPMM, they exhibit a particular property in the nGMM: they do not affect the global aggregate $(x,y)$. Thus, for any swap $\Delta x$ of $X$ sent to an AMM, any corresponding $\Delta y$ is compatible with maintaining a constant aggregate product.
Here, we adopt the natural exchange rate given by the ratio of reserves, $\frac{y}{x}$, so that $\Delta y_{nGMM}=\frac{y}{x}\Delta x$ ensures that each AMM preserves its value, computed at the current aggregate marginal prices, i.e.,
$(y_{i}-\Delta y_{nGMM})+\frac{y}{x}(x_{i}+\Delta x)=(y_{i}-\frac{y}{x}\Delta x)+\frac{y}{x}(x_{i}+\Delta x)=y_{i}+\frac{y}{x}x_{i}.$
Once the nGMM is defined for these inter-AMM swaps, the GMM can be defined as in Section 4, and the properties discussed there apply in the same manner to these swaps. Indeed, as we shall show later, these swaps are equivalent to other swaps with zero net trade implemented by an external agent. From now on, we assume that the GMM incorporates these specific pricing rules for inter-AMM trades.
Next, we introduce a modification of the GMM designed to guarantee traders an improvement over the CPMM.
Definition: The GMM algorithm with rebalancing is a global algorithm that applies the GMM algorithm to the resulting updated vector of reserves from executing the following algorithm iteratively until one of the conditions fails (assume wlog $\Delta x>0)$.
  1. (a)
    Check whether $l\in\arg\max\{x_{j}y_{j}\}_{j\in I}$.
  2. (b)
    Check whether (10) is met.
  3. (c)
    Check whether $\frac{\tilde{y}_{l}}{\tilde{x}_{l}}<r$, for the (updated) vector of reserves $(\tilde{x}_{l},\tilde{y}_{l})$.
  4. (d)
    Update reserves using a swap from AMM $l$ to $j\in\arg\max_{k\in I\setminus\{l\}}\frac{\tilde{y}_{k}}{\tilde{x}_{k}}$ of
    $\min\left\{\frac{r\tilde{x}_{j}-\tilde{y}_{j}}{2r},\frac{\tilde{y}_{l}-r\tilde{x}_{l}}{2r}\right\}$
    units of $X$.
  5. (e)
    Go to (c).
The GMM algorithm with rebalancing is a variation of the GMM algorithm that, before applying it, checks whether (10) holds. It is under this condition that Proposition 10 states that the AMM provides better terms of trade for the trader under (a) than (b). Specifically, the trader benefits from trading with the AMM that has the largest product of reserves (and thus the lowest slippage) after all arbitrage opportunities are exhausted and all AMMs have a reserve ratio equal to the global ratio.
In this case, the GMM algorithm with rebalancing redistributes swaps to the other AMMs until the AMM’s reserve ratio aligns with the global ratio. Consequently, the GMM with rebalancing offers marginal terms of trade at least as favorable as those in the CPMM with arbitrage. Additionally, slippage is lower since rebalancing increases reserves because the AMM captures a portion of the arbitrageurs’ profits.
Proposition 5.
Consider an initial distribution of liquidity $(\vec{x},\vec{y})$ and a swap $\Delta x>0$ of $X$ (wlog) that is optimally routed. Compare the following two scenarios:
  1. (a)
    All AMMs use the CPMM, and all strictly profitable arbitrage opportunities in the initial liquidity distribution are exploited by balanced arbitrage before the swap $\Delta x$ is executed.
  2. (b)
    All AMMs use the GMM algorithm with rebalancing.
(b) provides better terms of trade for the trader.
Proof.
Proposition 10 and the definition of GMM with rebalancing means that we can restrict to the case where (10) is satisfied. Suppose from now on that this is the case. In (a), arbitrage equalises the ratio of reserves of all AMMs. Solving that the product of reserves of each AMM remains constant and that each ratio of reserves must equal to $r$, one can use the CPMM formula to show that AMM $j$ would respond to the swap of $\Delta x$ sending back
$\frac{r}{1+\frac{\Delta x}{\sqrt{\frac{x_{j}y_{j}}{r}}}},$ (13)
so that the most profitable trade for the trader is the AMM with largest product of reserves, say $l$. To prove the proposition we shall show that the same AMM in (b), offers strictly better terms of trade. Since $l=\arg\max_{j\in I}x_{j}y_{j}$, the second inequality of (10) implies that
$\frac{r_{l}}{1+\frac{\Delta x}{x_{l}}}<\frac{r}{1+\frac{\Delta x}{\sqrt{\frac{x_{l}y_{l}}{r}}}}=\frac{r}{1+\sqrt{\frac{r}{r_{l}}\frac{\Delta x}{x_{l}}}},$ (14)
which implies that $\frac{y_{l}}{x_{l}}=r_{l}<r$, and thus there are other AMMs with ratios of reserves larger than $r$. The algorithm GMM with rebalancing iterates sending swaps from $l$ to each of other AMMs in step 4. At each iteration in which the updated ratio of reserves of $l$ and $j$, $\tilde{r}_{l}\equiv\frac{\tilde{y}_{l}}{\tilde{x}_{l}}$ and $\tilde{r}_{l}\equiv\frac{\tilde{y}_{j}}{\tilde{x}_{j}}$ respectively, satisfy $\tilde{r}_{l}<r<\tilde{r}_{j}$, there are two possibilities. Consider first that,
$\frac{r\tilde{x}_{l}-\tilde{y}_{l}}{2r}>\frac{\tilde{y}_{j}-r\tilde{x}_{j}}{2r},$ (15)
then the swap sent to $j$ is equal to $\tilde{\Delta}x=\frac{\tilde{y}_{j}-r\tilde{x}_{j}}{2r}$ units of $X$. Applying the GMM with rebalancing formula, the amount of $Y$ obtained with the swap by $l$ is equal to:
$\min\left\{\frac{\tilde{y}_{j}}{\tilde{x}_{j}+\tilde{\Delta}x},r\right\}\tilde{\Delta}x$ $=$ $\min\left\{\frac{\tilde{y}_{j}}{\tilde{x}_{j}+\frac{\tilde{y}_{j}-r\tilde{x}_{j}}{2r}},r\right\}\tilde{\Delta}x$
$=$ $\min\left\{\frac{\tilde{r}_{j}}{1+\frac{\tilde{r}_{j}-r}{2r}},r\right\}\tilde{\Delta}x$
$=$ $\min\left\{\frac{\tilde{r}_{j}}{r+r_{j}},1\right\}r\tilde{\Delta}x$
$=$ $r\tilde{\Delta}x.$
Thus, the swap sent to $j$ increases the ratio of reserves of $l$ to:
$\frac{\tilde{y}_{l}+r\tilde{\Delta}x}{\tilde{x}_{l}-\tilde{\Delta}x}=\frac{\tilde{y}_{l}+\frac{\tilde{y}_{j}-r\tilde{x}_{j}}{2}}{\tilde{x}_{l}-\frac{\tilde{y}_{j}-r\tilde{x}_{j}}{2r}}<\frac{\tilde{y}_{l}+\frac{r\tilde{x}_{l}-\tilde{y}_{l}}{2}}{\tilde{x}_{l}-\frac{r\tilde{x}_{l}-\tilde{y}_{l}}{2r}}=r.$
Thus, the algorithm goes back to 3. and then 4. again, repeating the process. Note that there are other AMMs with reserve ratios strictly larger than $r$ since the updated ratio of reserves of $l$ is strictly less than $r$ and $r$ is the aggregate ratio of reserves. Consider now the case:
$\frac{r\tilde{x}_{l}-\tilde{y}_{l}}{2r}\leq\frac{\tilde{y}_{j}-r\tilde{x}_{j}}{2r}.$ (16)
A similar argument as above implies that the ratio of reserves of $l$ increases to $r$ and then the algorithm stops at 3. once it revisits it again. Since this is the only way the iterative part of the algorithm stops. The amount of $Y$ that the algorithm returns to the trader must be the minimum between the nGMM quantity and the CPMM quantity that corresponds to the updated liquidity reserves. The former is equal to:
$\frac{y}{x+\Delta x}\Delta x=\frac{r}{1+\frac{\Delta x}{x}}\Delta x,$ (17)
and the latter to:
$\frac{\tilde{y}_{l}}{\tilde{x}_{l}+\Delta x}\Delta x=\frac{r}{1+\frac{\Delta}{\tilde{x}_{l}}}\Delta x=\frac{r}{1+\frac{\Delta}{x_{l}\left(\frac{r+r_{l}}{2r}\right)}}\Delta x,$ (18)
where $(\tilde{x}_{l},\tilde{y}_{l})$ denotes the final vector of liquidity reserves of $l$ after running the iterative part of the algorithm. Since this part only stops at the point in which $\frac{\tilde{y}_{l}}{\tilde{x}_{l}}=r$ and it satisfies $r(x_{l}-\tilde{x}_{l})=y_{l}-\tilde{x}_{l}$ we have that $(\tilde{x}_{l},\tilde{y}_{l})=\left(x_{l}\frac{r+r_{l}}{2r},x_{l}\frac{r+r_{l}}{2}\right)$. Since $x_{l}\frac{r+r_{l}}{2r}\leq x_{l}\leq x$, (18) is less than (17), and thus the amount of $\Delta y$ returned by the algorithm is equal to (18). To finish the proof we show that (18) is greater than (13) for $j=l$. This is equivalent to show that $\frac{r+r_{l}}{2r}\geq\sqrt{\frac{r_{l}}{r}}$, which can be easily checked. ∎
One concern, however, is that rebalancing may introduce perverse effects. We argue that this is not the case by showing that rebalancing produces the same outcome for the AMMs as a specific sequence of trades with zero net trade applied to the standard GMM. For simplicity, we illustrate this claim with an example, though the argument is general.
{exampleboxinner}
[frametitle=Toy example - Part 7: Rebalancing]
Suppose a set of two AMMs with initial vector of reserves
$(\vec{x},\vec{y})=(90ETH,210ETH,440000UST,760000UST),$
and a swap of 1 ETH sent by a trader to AMM (b). In this case, the GMM with balancing applies the iterative algorithm sending 10ETH to AMM 1 to get 40000UST in return ($\frac{1200000UST}{300ETH}*10ETH$). Then, the vector of reserves is updated to
$(100ETH,200ETH,400000UST,800000UST),$
and AMM 2 returns to the trader $3980,10UST\approx\frac{800000UST}{200+1}ETH*1ETH$ since it is a divergent order.
Consider next, the standard GMM and that the swap sent by the trader is preceded by the following two trades sent by an additional trader. The additional trader sends 10ETH to AMM1 to get 38709.67 UST in return ($\approx\frac{1200000UST}{300+10ETH}*10ETH$ since it is a convergent trade). This updates the vector of reserves to
$(100ETH,210ETH,436.129,03UST,760000UST),$
and then sends 38709.67 UST to AMM2 to get 10ETH in return ($\approx\frac{310ETH*38709.67UST}{1.161.290,32+38709.67UST}$ since it is again a convergent trade). After these two trades, the vector of reserves is updated to
$(100ETH,200ETH,400000UST,800000UST),$
and now when the original trade is submitted to AMM 2 it returns to the trader $3980,10UST\approx\frac{800000UST}{200+1}ETH*1ETH$ since it is a divergent order.
Thus, rebalancing neither enables additional manipulative strategies nor alters the cost of manipulation compared to using phony trades. The only caveat is that manipulation with phony trades requires additional asset holdings (e.g., 10 ETH in the example).
5.3 MEV Sandwich Attacks
Next, we show MEV sandwich are strictly less profitable when the AMM uses the GMM algorithm than when it uses the CPMM.
Definition: Given a swap of $\Delta x>0$ units of $X$ submitted to AMM $i$, a MEV sandwich attack of $\widehat{\Delta x}$ units of $X$ consists of a frontrunning swap and a backrunning swap submitted to AMM $i$ before and after $\Delta x$, respectively. The frontrunning swap sends $\widehat{\Delta x}$ units of $X$ and the backrunning swap sends the units of $Y$ obtained with the first swap.
A MEV sandwich may be profitable due to slippage: the first order buys $X$ more cheaply than the price at which the second order sells because there is an order in between that makes $X$ more scarce at the AMM. This is clearly the case when the AMM uses CPMM in which prices reflect the scarcity at the AMM but it is less so in the case of the GMM. In this later case, the first buy order is satisfied at a price that reflects the scarcity created by the order but the subsequent sell order is satisfied at a price that reflects the much smaller scarcity existing in the set of AMMs. The reason for this is that one would expect that AMMs should start from an equilibrium situation in which all (marginal) prices are the same (and thus local levels of scarcity) and thus the first buy order moves the AMM away from the other AMMs whereas the sell order moves the AMM towards the other AMMs. Next proposition formalises this argument.
Proposition 6.
The profit of an elementary MEV sandwich is strictly lower when the AMM uses GMM than when the AMM uses CPMM.
Proof.
By definition, the profit of ane elementary MEV sandwich is equal to $\widehat{\Delta^{F}x}_{s}-\widehat{\Delta x},$ for algorithm $s\in\{CPMM,GMM\}$, where:
$\widehat{\Delta}y_{CPMM}$ $=$ $\frac{y_{i}}{x_{i}+\widehat{\Delta}x}\widehat{\Delta}x,$
$\Delta y_{CPMM}$ $=$ $\frac{y_{i}-\widehat{\Delta}y_{CPMM}}{x_{i}+\widehat{\Delta}x+\Delta x}\Delta x,$
$\widehat{\Delta}^{F}x_{CPMM}$ $=$ $\frac{x_{i}+\widehat{\Delta}x+\Delta x}{y_{i}-\Delta y_{CPMM}}\widehat{\Delta}y_{CPMM},$
and:
$\widehat{\Delta}y_{GMM}$ $=$ $\min\left\{\frac{y_{i}}{x_{i}+\widehat{\Delta}x},\frac{y}{x+\widehat{\Delta}x}\right\}\widehat{\Delta}x$
$\Delta y_{GMM}$ $=$ $\min\left\{\frac{y_{i}-\widehat{\Delta}y_{GMM}}{x_{i}+\widehat{\Delta}x+\Delta x},\frac{y-\widehat{\Delta}y_{GMM}}{x+\widehat{\Delta}x+\Delta x}\right\}\Delta x,$
$\widehat{\Delta}^{F}x_{GMM}$ $=$ $\min\left\{\frac{x_{i}+\widehat{\Delta}x+\Delta x}{y_{i}-\Delta y_{GMM}},\frac{x+\widehat{\Delta}x+\Delta x}{y-\Delta y_{GMM}}\right\}\widehat{\Delta}y_{GMM}.$
Thus, the min implies that the profits are weakly less under the GMM than under the CPMM. To show that they are strictly less we only need to show that in one of the mins, the second term is strictly less then the first. For that, just note that if it were not the first one, the following chaing of implications implies that it has to be the third one:
$\frac{y_{i}}{x_{i}+\widehat{\Delta}x}$$\leq$ $\frac{y}{x+\widehat{\Delta}x}$
$\Rightarrow\frac{y_{i}}{x_{i}+\widehat{\Delta}x+\Delta x}$ $<$ $\frac{y}{x+\widehat{\Delta}x+\Delta x}$
$\Rightarrow\frac{x_{i}+\widehat{\Delta}x+\Delta x}{y_{i}}$ $>$ $\frac{x+\widehat{\Delta}x+\Delta x}{y}$
$\Rightarrow\frac{x_{i}+\widehat{\Delta}x+\Delta x}{y_{i}-\Delta y_{GMM}}$ $>$ $\frac{x+\widehat{\Delta}x+\Delta x}{y-\Delta y_{GMM}}.$
The first line says that in the first min, the left term is weakly less than the right one; the last line says that in the third min, the left term is strictly greater than the right one; the second implication is direct and the first one and the last one follow from the observation that $y_{i}(x+\beta)-y(x_{i}+\beta)$ it is strictly decreasing in $\beta$ for $y_{i}<y$. ∎
Next, we redo the MEV extraction example computed for CPMM in Example 2 in the case of GMM and we observe that the extracted MEV is so small that it can become negative.
{exampleboxinner}
[frametitle=Toy example - Part 7: MEV extraction in GMM]
Consider two AMMs using the GMM algoritm and with a liquidity pool containing 100 ETH and 400,000 UST each. The marginal price of ETH is then 4,000 UST/ETH. Suppose that a trader interested in converting 40,000 UST in ETH sends them to AMM 1. The MEV-extractor can perform a sandwich attack by sending 60,000 UST to AMM 1 just before the victim’s transaction is executed and to send the proceeds back to AMM 1 immediately after. This would work as follows. Note that along all trades the reserves of AMM 2 remain constant and equal to 100 ETH and 400,000 UST.
  • Frontrunning: MEV-extractor transaction 1 sends 60,000 UST to AMM 1 to get 13.0435 ETH in return since this order is divergent and thus priced as the CPMM. This order changes the vector of reserves of AMM 1 to 86,9565 ETH and 460.000 UST. Its marginal price of ETH increases to 5,290 UST/ETH.
  • Victim’s transaction: The victim sends 40,000 UST to AMM 1. This order is also divergent and thus priced as in the CPMM. The trader obtains in return 6.9565 ETH. The resulting reserves of AMM 1 are then 80 ETH and 500,000 UST. Nothing changes in AMM 2.
  • Backrunning: MEV-extractor transaction 2 sends 13.0435 ETH to AMM 1. In this case, the order is convergent and thus priced by the nGMM. Since the global reserves are 180 ETH and 900,000 UST, the MEV-extractor gets 60,811 UST. The MEV-extractor makes a net profit of just 811 UST, substantially less than the profit of 10,094 UST obtained in the CPMM.
We also provide a general analysis of the profits of an elementary MEV sandwich. In this case, we also need to specify the global reserves. To simplify the notation we assume that the ratio of global reserves $\frac{t}{x}$ is equal to the ratio of reserves of AMM i $\frac{y_{i}}{x_{i}}$.
Proposition 7.
The profit of an elementary MEV sandwich $\hat{\Delta}x$ to a trader order $\Delta x$ to an AMM i with liquidity reserves $(x_{i},rx_{i})$ and global reserves $(x,rx)$ running the GMM algorithm is equal to:
$\left(\frac{\left(1+\frac{\hat{\Delta}x}{x}+\frac{\Delta x}{x}\right)\left(1+\frac{\hat{\Delta}x}{x_{i}}+\frac{\Delta x}{x_{i}}\right)}{\left(1+\frac{\hat{\Delta}x}{x_{i}}+\frac{\Delta x}{x_{i}}\right)\left(1+\frac{\hat{\Delta}x}{x_{i}}\right)-\frac{\Delta x}{x}}-1\right)\hat{\Delta}x$ (19)
Proof.
As in the proof of Proposition 6, we can compute the profits of MEV extractor as $\hat{\Delta}^{F}x-\hat{\Delta}x$ where:
$\begin{array}[]{rcl}\widehat{\Delta}y=&\min\left\{\frac{r}{1+\frac{\widehat{\Delta}x}{x_{i}}},\frac{r}{1+\frac{\widehat{\Delta}x}{x}}\right\}\widehat{\Delta}x&=\frac{r}{1+\frac{\widehat{\Delta}x}{x_{i}}}\widehat{\Delta}x,\\ \Delta y=&\min\left\{\frac{r-\frac{\widehat{\Delta}y}{x_{i}}}{1+\frac{\widehat{\Delta}x}{x_{i}}+\frac{\Delta x}{x_{i}}},\frac{r-\frac{\widehat{\Delta}y}{x}}{1+\frac{\widehat{\Delta}x}{x}+\frac{\Delta x}{x}}\right\}\Delta x&=\frac{r-\frac{\widehat{\Delta}y}{x_{i}}}{1+\frac{\widehat{\Delta}x}{x_{i}}+\frac{\Delta x}{x_{i}}}\Delta x,\\ \widehat{\Delta}^{F}x=&\min\left\{\frac{1+\frac{\widehat{\Delta}x}{x_{i}}+\frac{\Delta x}{x_{i}}}{r-\frac{\Delta y}{x_{i}}},\frac{1+\frac{\widehat{\Delta}x}{x}+\frac{\Delta x}{x}}{r-\frac{\Delta y}{x}}\right\}\widehat{\Delta}y&=\frac{1+\frac{\widehat{\Delta}x}{x}+\frac{\Delta x}{x}}{r-\frac{\Delta y}{x}}\widehat{\Delta}y.\end{array}$
We can derive the proposition from the recursive substitution of $\Delta y$ and $\widehat{\Delta}y$ and some straightforward simplifications. ∎
Figure 6 illustrates the proposition for different values of $\hat{\Delta}x$ and compare the results with MEV profits obtained in the CPMM case. The GMM reduces substantially the profitability of MEV attacks and even render them unprofitable.
Figure 6: The graph plots the profits of elementary MEV sandwich manipulations $\hat{\Delta}x$ for $x_{i}=400,000$, $x=800.000$ and $\Delta x=40,000$.
5.4 Impermanent Losses
In the case of impermanent losses for the GMM we need to keep track how traders route their orders as the price changes from $r_{0}$ to $r$ since unlike in the CPMM, this affects how the value of the reserves of the AMM evolve.
We adopt a simple setting that we refer as the ideal benchmark. In the ideal benchmark, we assume that there are two AMMs only $1$ and $2$, with respectively reserves $(x_{1},y_{1})$ and $(x_{2},y_{2})$. As in the case of the CPMM, we also assume that the initial ratio of reserves of both AMMs is equal to the initial price $r_{0}$, i.e. $\frac{y_{1}}{x_{1}}=\frac{y_{2}}{x_{2}}=r$ and we model a change in market sentiment assuming an insider trader that submit trades to the AMM to maximize the value of its net final position at a the new price $r^{\prime}$.
Note that the impermanent losses under the ideal benchmark are an upper bound as adding additional trades can only increase the value of the final liquidity reserves of the AMMs.
Finally, we assume without loss of generality that $\alpha\equiv\frac{x_{2}}{x_{1}+x_{2}}\leq 0.5$, which under the above assumptions means that AMM 1 is the AMM with largest initial market value $rx_{i}+y_{i}$.
Proposition 8.
In the ideal benchmark, the impermanent losses of AMM 1 are the same as in the case of the CPMM, whereas the impermament losses of AMM 2 are equal to:
$IL_{2}=1-2\frac{\sqrt{\left(\sqrt{\frac{r^{\prime}}{r}}+\frac{1-\alpha}{\alpha}\right)\left(\sqrt{\frac{r}{r^{\prime}}}+\frac{1-\alpha}{\alpha}\right)}-\frac{1-\alpha}{\alpha}}{\sqrt{\frac{r^{\prime}}{r}}+\sqrt{\frac{r}{r^{\prime}}}}.$ (20)
Proof.
Assume $r^{\prime}<r$. The other case is symmetric. Since AMM 1 slippage is lower but marginal prices are the same, it is optimal for the insider trader to trade with AMM 1 first. This first order is divergent and thus priced by CPMM prices, so the quantity $\Delta x^{*}_{1}$ of $X$ sent to AMM 1 solves:
$\max_{\Delta x_{1}}\left(\frac{r}{1+\frac{\Delta x_{1}}{x_{1}}}-r^{\prime}\right)\Delta x_{1},$
this is $\Delta x^{*}_{1}=x_{1}\left(\sqrt{\frac{r}{r^{\prime}}}-1\right)$, and thus,
$\Delta y^{*}_{1}\equiv\frac{r}{1+\frac{\Delta x^{*}_{1}}{x_{1}}}\Delta x_{1}^{*},$
is the quantity of $Y$ received by the insider trader. Note that the optimal value $\Delta x_{1}^{*}$ is such that the updated ratio of reserves of AMM 1 is equal to the new price $r$:
$\frac{y_{1}-\Delta y^{*}_{1}}{x_{1}+\Delta x^{*}_{1}}=\frac{r-\frac{r}{1+\frac{\Delta x^{*}_{1}}{x_{1}}}\frac{\Delta x^{*}_{1}}{x_{1}}}{1+\frac{\Delta x^{*}_{1}}{x_{1}}}=\frac{r}{\left(1+\frac{\Delta x^{*}_{1}}{x_{1}}\right)^{2}}=r^{\prime},$ (21)
which means that $(x_{1}+\Delta x^{*}_{1})(y_{1}-\Delta y^{*}_{1})=r^{\prime}(x_{1}+\Delta x^{*}_{1})^{2}$, and thus that,
$x_{1}=\sqrt{\frac{r^{\prime}}{r}}\left(x_{1}+\Delta x^{*}_{1}\right),$ (22)
since $(x_{1}+\Delta x^{*}_{1})(y_{1}-\Delta y^{*}_{1})=x_{1}y_{1}=rx_{1}^{2}$.
The second optimal order for the insider trader is to send to AMM 2, the convergent order $\Delta x^{*}_{2}$ of $X$ that solves:
$\max_{\Delta x_{2}}\left(\frac{\hat{r}}{1+\frac{\Delta x_{2}}{x^{U}}}-r^{\prime}\right)\Delta x_{2},$
and returns an amount of $Y$ that we denote by $\Delta y^{*}_{2}$. By a similar argument as for $\Delta x_{1}^{*}$:
$\frac{x_{2}+\Delta x_{2}^{*}+x_{1}+\Delta x_{1}^{*}}{y_{2}-\Delta y_{2}^{*}+y_{1}-\Delta y_{1}^{*}}=r^{\prime},$ (23)
which together with (21) means that:
$\frac{x_{2}+\Delta x_{2}^{*}}{y_{2}-\Delta y_{2}^{*}}=r^{\prime}.$ (24)
Finally, the nGMM also means that:
$(x_{2}+\Delta x_{2}^{*}+x_{1}+\Delta x_{1}^{*})(y_{2}-\Delta y_{2}^{*}+y_{1}-\Delta y_{1}^{*})=(x_{2}+x_{1}+\Delta x_{1}^{*})(y_{2}+y_{1}-\Delta y_{1}^{*}),$
which can be transformed into:
$r^{\prime}\left(x_{2}+\Delta x_{2}^{*}+\frac{1-\alpha}{\alpha}\sqrt{\frac{r}{r^{\prime}}}x_{2}\right)^{2}=rx_{2}^{2}\left(1+\frac{\alpha}{1-\alpha}\sqrt{\frac{r}{r^{\prime}}}\right)\left(1+\frac{\alpha}{1-\alpha}\sqrt{\frac{r^{\prime}}{r}}\right),$
using $x_{1}=\frac{1-\alpha}{\alpha}x_{2}$, $y_{2}=rx_{2}$, and (21)-(24). Thus, solving for $x_{2}+\Delta x_{2}^{*}$ gives us:
$x_{2}+\Delta x_{2}^{*}=x_{2}\sqrt{\frac{r}{r^{\prime}}}\left(\sqrt{\left(1+\frac{\alpha}{1-\alpha}\sqrt{\frac{r}{r^{\prime}}}\right)\left(1+\frac{\alpha}{1-\alpha}\sqrt{\frac{r^{\prime}}{r}}\right)}-\frac{1-\alpha}{\alpha}\right).$
We use this expression to compute the value of AMM 2 at the final prices and reserves.
$V_{2}$$\equiv$ $r^{\prime}(x_{2}+\Delta x^{*}_{2})+(y_{2}-\Delta y^{*}_{2})$
$=$ $2r^{\prime}(x_{2}+\Delta x^{*}_{2})$
$=$ $2r^{\prime}x_{2}\sqrt{\frac{r}{r^{\prime}}}\left(\sqrt{\left(1+\frac{\alpha}{1-\alpha}\sqrt{\frac{r}{r^{\prime}}}\right)\left(1+\frac{\alpha}{1-\alpha}\sqrt{\frac{r^{\prime}}{r}}\right)}-\frac{1-\alpha}{\alpha}\right),$
whereas the value of the initial reserves of AMM 2 at the new price is:
$V_{2}^{0}=r^{\prime}x_{2}+y_{2}=x_{2}(r^{\prime}+r).$
which together with the definition of impermanent losses as $IL_{2}=1-\frac{V_{2}}{V^{0}_{2}}$ implies the proposition. ∎
Figure 7 illustrates Proposition 20 and compares the impermananent losses in the GMM with respect to the CPMM.
Figure 7: The graph plots the impermanent losses as a function of the price variation $\frac{r}{r_{o}}$ for different values of $\alpha$ in the ideal benchmark using Equation (20).
The next example illustrate this result.
{exampleboxinner}
[frametitle=Toy example - Part 8: Impermanent losses in GMM] Consider two GMMs with liquidity pools containing 100 ETH and 400,00 UST. The marginal price is then 4,000 UST/ETH.
Similarly than in Example 3, suppose that one or more traders submit transactions for buying 30 ETH on each GMM.
We can compute impermanent loss IL) as the difference between the value of the GMM’s original reserves valued at the current price and the current reserves valued at the current price.
To make a simple example, we assume that a first order for 30 ETH arrives to GMM1 and a second order for 30 ETH arrives to GMM2. The final reserves for GMM1 are the same as in one of the CPMMs in Example 3, namely 70 ETH and 571,429 UST.
The final reserves of GMM2 resulting from the second trade of 30 ETH can be computed using the convergent formula for 28 ETH, at point which the relative prices are equalized for the two GMMs and the divergent formula for the remaining 2 ETH.
Overall, this implies that the impermanent loss for GMM1 and GMM2 are equal to 129,023 UST, which is 17,899 UST less than in Example 3.
The next section provides simulations using real life transactions and provides a quantitative assessment of how large this reduction of impermanent losses is.
6 Empirical Evaluation
We proceed with evaluating the theoretical results with simulations based on real trades. We consider transactions in two AMMs in the Ethereum network, Uniswap V2 and Sushiswap. [^8] We chose Uniswap V2 instead of V3 due to the similar fee structure that it has compared to Sushiswap for ease of exposition. Moreover, we assume that all fees return to liquidity pools for both AMMs and ignore the 0.25% that is reserved in the treasury for Sushiswap The datasets are used to illustrate each of the metrics discussed in previous sections: arbitrage profits, profist of MEV sandwich attacks and impermament losses. We flag arbitrage and sandwich operations using the identification of the ZeroMev database. [^9] https://zeromev.org/
We start by providing an assessment of the benefits of replacing the CPMM with the GMM in terms of the cost of arbitrage. We consider 3 periods for one of the most traded pairs USDT/WETH, two periods with high price variation and one with low. Moreover, we consider three other less traded pairs: ALCX/WETH, BLOCKS/WETH and VSP/WETH for periods that they display large arbitrage profits. We were not systematic on the section of pairs or periods beyond what we already explained. Table 1 presents the description of the samples and their characteristics. Table 6 confirms that routing is not far away from the optimal routing assumption that we introduced in Subsection 5.1. Except for Blocks and VSP, the cost of suboptimally routing orders only means a cost in the range of a few basis points.
Pairs WETH/ Ethereum Blocks Timeframe Transactions Description
USDT (1) 11,565,020 - 11,640,849 Aug 2021-12 days 114,974 tx High volatility
USDT (2) 12,390,294 - 12,467,913 May 2021-12 days 153,944 tx High volatility
USDT (3) 13,317,339 - 13,406,729 Oct 2021-12 days 59,156 tx Low volatility
ALCX 11,948,971 - 12,050,004 Jul 2021-20 days 21,481 tx Low Volatility / Unequal pools
BLOCKS 14,356,780 - 14,497,019 Mar 2022-20 days 4,297 tx High Volatility / Equal Pools
VSP 13,528,195 - 13,917,471 Nov 2021-20 days 2,681 tx High Volatility / Equal Pools
Table 1: Samples
Pairs WETH/ Traders losses compared to optimal routing / Volume of Trade
USDT (1) 0.03 %
USDT (2) 0.07%
USDT (3) 0.016 %
ALCX 0.15%
BLOCKS 1.7 %
VSP 2.3%
To estimate the potential gains of replacing the CPMM with GMM with respect to the cost of arbitrage we use the arbitrage profits as discussed in Subsection 5.1. We provide this measure of the cost of arbitrage with respect to the revenue fees of the AMMs which may be interpreted as the cost of liquidity provision by the AMMs. Our results show that the gains for a traded pair are modest ranging from 0.23% to 5.82%. For the other less traded pairs, the gains are significant which suggests that the GMM can be particularly effective for less traded pairs.
Pairs WETH/ Fees Arbitrage Arbitrage/Fees Ratio
USDT (1) 2,659 WETH 154.8 WETH 0.0582
USDT (2) 4,187 WETH 138.19 WETH 0.0330
USDT (3) 719.3 WETH 1.67 WETH 0.0023
ALCX 466.25 WETH 192.25 WETH 0.4124
BLOCKS 27.4 WETH 190.93 WETH 6.9701
VSP 21.18 WETH) 53.7 WETH 2.5368
Table 2: Arbitrage Profits and Fees
Lastly, we provide an estimate on the size of arbitrage based on a larger sample of three months (blocks 18251371 to 18908876, 01/10/23 to 31/12/23) for Uniswap V2. For this period, a total of 71.6 million USD was collected in fees while 3.4 million was extracted from arbitrageurs. The problem seems to be much more prevalent in smaller/less traded pairs. Specifically, out of 1,833 pairs that experienced at least one arbitrage operation, $20\%$ exhibited Arbitrage Profits to Fees ratios greater than $10\%$. These pairs represented approximately $7.45\%$ of the total AMMs fees collected in the period, an approximate meausure of the relative volume traded. Specifically these AMMs collected 5.33 million USD in fees while they experienced 2.56 million USD of arbitrage profits. This confirms the observation that the GMM may be particularly useful for less traded pairs that tend to have large arbitrage costs, measured by arbitrage profits, relative to liquidity provision costs, measured by the AMMs’ revenue from fees.
Next we proceed with assessing assess the effectiveness of the GMM algorithm in reducing the profitability of sandwich operations. To do so we analyze all sandwich operations that occurred for the most traded pairs with WETH for the period 2021 and 2023. For the simulations, we use the reserves at the beginning of each block where a sandwich operation appears and simulate the entire block of transactions affecting the AMMs of interest, Uniswap v2 and Sushiswap. Then, we compare the profit of the sandwich operation under CPMM and GMM. For the counterpart AMM (Sushiswap), when simulating the GMM, we use its reserves at the start of the same block but keep them constant, disregarding trades that may have occurred in parallel with the sandwich operation. To the extend that the excluded trades might increase the reserves of the AMMs, and that larger reserves make sandwich operations less profitable, our measure of sandwich profits for the GMM is an upper bound. Some sandwich attacks identified by zeromev seem to be substantially more complex than the elementary MEV sandwichs we describe in Section 5. If this is the case, we measure the profits as in Section 5 replacing the frontrun operation operation in the data with the corresponding backrun operation that makes the attack an elementary MEV sandwich. [^10] Measuring the profits of general MEV sandwich positions is an open question in the literature as it is usually the case in reality that the MEV sandwich attack is combined with other operations to economize gas usage. We find a more general estimate of MEV profist is beyond the scope of this paper. If an operation after this adjustment becomes unprofitable for the CPMM, we discard it. Our simulations suggest that this exclusion does not introduce significant bias, as our simulated MEV operators’ profits explain approximately $80\%$ of the profits reported by zeromev. Our results, presented in Table 3, show that in all cases, the GMM design significantly reduces the profitability of sandwich operations and, in many cases, renders them unprofitable, which suggest a substantial reduction of MEV attacks.
Pairs WETH/ Profitable Sandwich Attacks (CPMM) Profitable Sandwich Attacks (GMM) Sandwich Profits (CPMM) Sandwich Profits (GMM)
USDT (Uni) 446 177 \$988,254 \$260,000
USDC (Uni) 240 90 \$853,872 \$140,000
DAI (Uni) 443 94 \$922,269 \$84,311
WBTC (Uni) 123 18 4 WETH / 20 BTC 0.1 WETH / 0.78 BTC
LINK (Uni) 464 200 8,949 LINK / 85 WETH 1,479 LINK / 9.86 WETH
LDO (Sushi) 1,152 1,114 139,336 LDO / 260 WETH 132,183 LDO / 245 WETH
CRV (Sushi) 659 473 140,345 CRV / 197 WETH 21,102 CRV / 90 WETH
Table 3: Simulations for Sandwich operations
Finally, similarly to arbitrage we present our findings on the size of sandwich for the same longer period considered for arbitrage operations. Again, we measure the cost of sandwich attacks for traders with the profitability for the attackers, and compare it to the cost of liquidity provision measured by the revenue of AMMs from fees. In that period and for all the AMMs that exhibited at least one sandwich operation, the cost of sandwich attacks for traders is 90 million USD, which is close to the cost of liquidity provision of 94.7 million USD. Furthermore, out of 11,188 pairs that exhibited a sandwich operation, 94% of them had a ratio of trader’s sandwich cost to liquidity cost of 10% or more. Those AMMs collected a total of 36 million USD in fees while sandwitch operators profited 89.8 million USD.
Finally, we conclude our analysis by computing the impermament losses using real data and compare the CPMM with a simulated counterfactual corresponding to the GMM. We use data from three pairs. The first pair USDT/WETH is a well traded pair where both Sushiswap and Uniswap have relatively deep and equal pools. The second pair ALCX/WETH was chosen to demonstrate how the GMM performs when the reserves of the two AMMs are highly concentrated (Sushiswap pocesses more than $99\%$ of the total reserves for the period). Lastly, VSP/WETH represents a pair with relatively shallow but equal pools.
We add a simulated trend to the original data. We consider two cases corresponding to a price increase of $50\%$ ($r=1.5$) and $100\%$ ($r=2$) for the whole period. To make the prices fit this trend, we take the real transactions and increase orders for WETH and decrease orders for the counterpart currency for each consecutive day to fit the intended price increase for the whole period. Specifically, for the first day we sample the average of demand for WETH, we increase it by 2% and draw new trades, for the second day we increase the average by 4% and so on. Table 4 describes the results.
Pairs WETH/ Reserve Distribution Initial Price of WETH Time (Block #) Trend I.L. Difference / I.L. Total
USDT 48 - 52% 3,484 Aug 2021 (Block: 13,317,339) r=1.5 3.5%
USDT 48 - 52% 3,484 Aug 2021 (Block: 13,317,339) r=2 5.2%
ALCX 0.01 - 99.9% 1.83 Jul 2021 (Block: 11,948,971) r=1.5 0.2%
ALCX 0.01 - 99.9% 1.83 Jul 2021 (Block: 11,948,971) r=2 0.3%
VSP 62 - 38% 656 Sep 2021 (Block: 13,528,471) r=1.5 4.3%
VSP 62 - 38% 656 Sep 2021 (Block: 13,528,471) r=2 6.5%
Table 4: Impermanent Loss Comparison
7 Conclusion
In this paper, we have analyzed how to improve an existing pricing algorithm used in automated market makers by leveraging economic analysis. Our focus has been on three key metrics: arbitrage costs, MEV sandwich attacks, and impermanent losses.
However, our investigation leaves aside some important aspects that warrant further research. The first is the performance of our algorithm, the GMM, when adopted by only a subset of market participants. Understanding this is crucial to determining whether there are incentives to transition toward a market structure in which all AMMs adopt the GMM. Additionally, it is relevant to assess whether the GMM benefits smaller players by improving their competitiveness or, conversely, reinforces the market power of larger players. Furthermore, partial adoption of the GMM introduces new trade-offs, as arbitrage opportunities are not fully eliminated unless the transition is complete.
The second aspect is the role of fees. Existing AMM algorithms, such as the CPMM, use fees to compensate liquidity providers and mitigate impermanent loss. The GMM, however, increases the reserves within the AMM without requiring fees. Additionally, by reducing impermanent losses, the GMM allows for lower fees without compromising the compensation of liquidity providers. A deeper analysis of these factors would provide a more comprehensive understanding of how the efficiency gains from the GMM are distributed between traders and liquidity providers.
Finally, our work introduces a new field of economic analysis: the design of smart contracts. This approach is related to the methodology used by Sönmez (2024) in designing market mechanisms for kidney exchange and school allocation, among other applications. Instead of optimizing based on traditional criteria such as revenue or surplus—subject to constraints like incentive compatibility—we adapt an existing practical implementation, the CPMM algorithm, to incorporate a meaningful improvement: better information gathering. This adaptation is constrained by desirable properties, such as ensuring that neither the AMM nor traders are worse off in a well-defined sense. We believe that this market design perspective on smart contract development can lead to further valuable applications in decentralized finance and beyond.
In this paper, we have investigated with the help of the tools of economic analysis how to improve an existing pricing algorithm used in automated market makers in three relevant metrics: cost of arbitrage, MEV sandwich attacks and impermanent losses.
In our investigation, we have left aside a few important elements that are worth investigating. The first one is the performance of our algorithm, the GMM, when it is adopted by only a subset of the market participants. In particular, this is relevant to understand whether there are incentives to transit to the market structure where all AMMs use the GMM algorithm. It is also relevant to understand whether the GMM helps smaller players to compete with larger players or it makes it easier to big players to defend their existing position. Finally, the performance of the market with partial implementation of the GMM also implies novel trade offs as partial implementation does not fully eliminate arbitrage opportunities.
The second one is the role of fees. Existing algorithms like the CPMM use the fees to retribute liquidity providers and compensate for impermanent losses. The GMM, however, increases the reserves deposited in the AMM without any need of fees. Besides, lower impermanent losses make it possible to reduce fees without compromising the compensation of liquidity providers. Taken these aspects into account allows for a more comprehensive approach on how the efficiency gains of the GMM are shared between the traders and liquidity providers.
Finally, our paper proposes a new field in which to apply economic analysis: the design of smart contracts. Our approach is also novel though related to the approached adopted by Sönmez (2024) to the design of mechanisms for kidney exchange or school allocation, among others. Instead of optimazing with respect with respect to a standard criterium like revenue or surplus, subject to some constraints, like incentive compatibility, we study how to adapt a practical implementation, the CPMM algorithm, to include a potential improvement, better information gathering, subject to some desirable properties, like neither the AMM nor trades are worse off in some particular sense. We believe this approach should also give rise to other fruitful applications in the design of smart contracts.